Hh Electrical devices, housed ih a metl 2h 슨re-produced lag A Process in whicl 307. are electrically defecttve) and 40%. have dented cans (c) prion to assembly Ascu me they are rejected foreihhed reason a) What is the prolbability hat a. device chosen et rendom will be a rejech? Ans, o.58 らsuppose Ithe electrical defects fall into two catego ries correepo nding to an c r,too high (E2)·Te P(퇴=0.1 and P(es) output voltage which is either too low CE) 0.2 (and P (c)=04 as loefore), what (s reject for either or both Ei and C? Ans.0.4% C r all? Ans, 0,5% as th a) the probabilty ot tinding a reject if we for either E1, 2,
Answer #1

a)probability that a device rejected =1-P(it has none of defect) =1-(1-0.3)*(1-0.4)=0.58

b)P(reject for either E1 or C1 or both) =P(E1)+P(C)-P(E1nC) =P(E1)+P(C)-P(E1)*P(C)=0.1+0.4-0.1*0.4 =0.46

c)

P(reject for either E1;E2 or C or all)= P(E1 U E2 UC)

=P(E1)+P(E2)+P(C)-P(E1)P(E2)-P(E1)P(C)-P(E2)P(C)+P(E1)P(E2)P(C)=

=0.1+0.2+0.4-0.1*0.2-0.2*0.4-0.1*0.4+0.1*0.2*0.4=0.568